The following selected account balances appeared on the financial statements of the Washington Company:    

Accounts Receivable, Jan. 1                                                 $13,000

Accounts Receivable, Dec. 31                                                  9,000

Accounts Payable, Jan 1                                                          4,000

Accounts Payable Dec. 31                                                        7,000

Merchandise Inventory, Jan 1                                                  10,000

Merchandise Inventory, Dec. 31                                               15,000

Sales                                                                                      56,000

Cost of Goods Sold                                                                   31,000

Question:

Cash Collections from customers are?

Cash paid to suppliers is?
Answer #1

Cash Collections from customers = 13000 + 56000-9000 = $60000

Purchase = 15000 + 31000 -10000 = $36000

Cash paid to suppliers = 4000 + 36000-7000 =$33000